Chapter 09 Homework Saved | 18 Required information The following information applies to the questions displayed below.] PartRequired 1A Required 1B Prepare journal entries to record above transactions and adjustments for 2018. View transaction listChapter 09 Homework A Saved 19 Required information [The following information applies to the questions displayed below.] ParChapter 09 Homework i Saved Required information [The following information applies to the questions displayed below.) Part 3Chapter 09 Homework i Saved 21 Required information Part 4 of 5 The following information applies to the questions displayedChapter 09 Homework Saved 22 Required information [The following information applies to the questions displayed below.] Part
Answer #1
1.1
Date General Journal Debit Credit
11-Nov Cash 4,800
Sales 4,800
11-Nov Cost of goods sold 900 =60*15
Merchandise inventory 900
30-Nov Warranty expense 288 =4800*6%
Estimated warranty liability 288
9-Dec Estimated warranty liability 180 =12*15
Merchandise inventory 180
16-Dec Cash 14,400
Sales 14,400
16-Dec Cost of goods sold 2,700 =180*15
Merchandise inventory 2,700
29-Dec Estimated warranty liability 360 =24*15
Merchandise inventory 360
31-Dec Warranty expense 864 =14400*6%
Estimated warranty liability 864
1.2
Date General Journal Debit Credit
5-Jan Cash 9,600
Sales 9,600
5-Jan Cost of goods sold 1,800 =120*15
Merchandise inventory 1,800
17-Jan Estimated warranty liability 435 =29*15
Merchandise inventory 435
31-Jan Warranty expense 576 =9600*6%
Estimated warranty liability 576
2
Warranty expense for November 2017 288 =4800*6%
Warranty expense for December 2017 864 =14400*6%
3 Warranty expense 576 =9600*6%
4
Warranty expense for November 2017 288
Warranty expense for December 2017 864
Less: Cost of replacements -540 =(12+24)*15
Estimated Warranty Liability balance 612
5 Estimated Warranty Liability balance, beginning 612
Warranty expense for January 2018 576
Less: Cost of replacements -435 =29*15
Estimated Warranty Liability balance 753